OpenStudy (anonymous):

A 21-foot ladder is leaning against a building. The base of the ladder is 7 feet from the base of the building.

OpenStudy (anonymous):

i got 18 degrees

OpenStudy (anonymous):

what is to find?

OpenStudy (anonymous):

most of the mutiple choice answers seem unlikely

OpenStudy (anonymous):

To the nearest degree, what is the measure of the angle that the ladder makes with the ground?

OpenStudy (anonymous):



OpenStudy (anonymous):

we figure out the cos right

OpenStudy (anonymous):

adjacent/ base

OpenStudy (anonymous):

cos theta = 3

OpenStudy (anonymous):

ya \[\theta = \cos^{-1} 3\]

OpenStudy (anonymous):

dang it says error on my calculator

OpenStudy (anonymous):

it cannot be calculated cos lies between 0 and 1

OpenStudy (anonymous):

so would i have to do 21/90=7/x

OpenStudy (anonymous):

which grade are you in?

OpenStudy (anonymous):

12 i messed up in that last post my bad

OpenStudy (anonymous):

you need to write the answer as theta = cos^-1 3

OpenStudy (anonymous):

we have to fin the answers in degrees from the ground

OpenStudy (anonymous):

 you need this angle right?

OpenStudy (anonymous):

yeast we have the nearest degree for a 21foot ladder with a base of 7

OpenStudy (anonymous):

cos-1(3) in the calculator says its error

OpenStudy (anonymous):

it cannot be calculated dude! dont you have learnt that cos lies between 0-1?

OpenStudy (anonymous):

i know that why iam stuck but i only have like 4 multiple choice answers

OpenStudy (anonymous):

give me that options

OpenStudy (anonymous):

18,71,72,67

OpenStudy (anonymous):

i think its 18 its my guess

OpenStudy (anonymous):

rewrite the whole question as it is again

OpenStudy (anonymous):

To the nearest degree, what is the measure of the angle that the ladder makes with the ground?

Directrix (directrix):

The figure @Timtime and @miteshchvm you have is correct. The suggested answer of 18 is not correct. cos(x) = 7/21 where x is the angle in degrees that the bottom of the ladder makes with the ground. Cranking that out gives a degree measure that can be entered into the calculator as cos-1(7/21) or cos-1(1/3). The answer is one of those posted above as options as you will see when you do the calculation. Be sure that your calculator is set to degrees and not radians or grads.